who has more mass?
N=14    H=1    O=16     C=12   S=32

a. 1 mol of N

b. 15 moles of H

c. 0.5 mol of S

e. 1 mol of C

d. 0.5 mol of O



1
Expert's answer
2021-01-15T06:49:57-0500

a ) 1 mol of N have " 14 g " .

b ) 15 mol of H have " 15 g " .

c ) 0.5 mol of S have 0.5 × 32 = 16 g .

d ) 1 mol of C have 1 × 12 = 12 g .

e ) 0.5 mol of O have 0.5 × 16 = 8 g .


Answer Should be " C " .

0.5 mol of S is the Answer .